You are scheduled to receive 14000 dollars in 2 years. When you receive it, you will invest it for 8 more years at 9.5 percent per year. How much you will have in 10 years?
Answer #1
Value of money in 10 years = P2 * (1+i)^n Where,
= $      14,000.00 * (1+0.095)^8 P2 = $ 14,000.00
= $      14,000.00 * 2.066869009 i = 9.50%
= $      28,936.17 n = 8